OpenStudy (anonymous):

How would you simplify: (-2+2i)^2

OpenStudy (cornzyblack):

thats [(-2)^2 + (2i)^2] = 4 +(-4) = 0

OpenStudy (anonymous):

\[(-2+2i)^2=(-2+2i)(-2+2i)=(4-4)+(-4-4)i=-8i\]

OpenStudy (anonymous):

there is no distributive law for exponents

OpenStudy (anonymous):

yeah.. .

OpenStudy (anonymous):

Where do all the 4's come from?